Consider the following system of differential equations representing a prey and predator
population model

dx/dt=x square -y

dy/dt= x+y
i) Identify all the real critical points of the system
ii) Obtain the type and stability of these critical points.
1
Expert's answer
2020-04-13T11:00:29-0400

QUESTION(i)

To find critical points, we must solve the system




"\left\{\begin{array}{l}
\displaystyle\frac{dx}{dt}=x^2-y\\[0.3cm]
\displaystyle\frac{dy}{dt}=x+y\end{array}\right.\longrightarrow
\left\{\begin{array}{l}
u(x,y)=x^2-y\\[0.3cm]
v(x,y)=x+y\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
x^2-y=0\\[0.3cm]
x+y=0
\end{array}\right.\longrightarrow
\left\{\begin{array}{l}
y=x^2\\[0.3cm]
x+x^2=0
\end{array}\right.\longrightarrow\left\{\begin{array}{l}
y=x^2\\[0.3cm]
x(x+1)=0
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=x^2\\[0.3cm]
x=0\quad\text{or}\quad x=-1
\end{array}\right.\longrightarrow\\[0.3cm]
\boxed{A(0,0)\,\,\,\text{and}\,\,\,B(-1,1)-\text{the real critical points of the system
}}"

QUESTION(ii)

Next we find the Jacobian matrix of 



"J(x,y)=\left(\begin{array}{cc}
\displaystyle\frac{\partial u}{\partial x} &\displaystyle\frac{\partial u}{\partial y}\\[0.3cm]
\displaystyle\frac{\partial v}{\partial x} &\displaystyle\frac{\partial v}{\partial y} 
\end{array}\right)=
\left(\begin{array}{cc}
2x&-1\\[0.3cm]
1 &1 \end{array}\right)"

Hints on remaining parts

  1. Find the eigenvalues of the Jacobian at each critical point.
  2. Classify the stability using the following.


In our case,


"A(0,0) : J(0,0)=
\left(\begin{array}{cc}
0&-1\\[0.3cm]
1 &1 \end{array}\right)\\[0.3cm]
\left|\begin{array}{cc}
0-\lambda&-1\\[0.3cm]
1 &1-\lambda \end{array}\right|=0-\text{characteristic equation}\\[0.3cm]
-\lambda(1-\lambda)-1\cdot(-1)=-\lambda+\lambda^2+1=0\\[0.3cm]
D=(-1)^2-4\cdot1\cdot1=1-4=-3\to\sqrt{D}=i\sqrt{3}\\[0.3cm]
\left[\begin{array}{l}
\lambda_1=\displaystyle\frac{1+i\sqrt{3}}{2}\\[0.3cm]
\lambda_2=\displaystyle\frac{1-i\sqrt{3}}{2}\end{array}\right.\longrightarrow Re(\lambda_1,\lambda_2)=\frac{1}{2}>0"



Conclusion,

The eigenvalues are complex conjugates. Real parts positive - An Unstable Spiral : All trajectories in the neighborhood of the fixed point spiral away from the fixed point with ever increasing radius.



"B(-1,1) : J(-1,1)=
\left(\begin{array}{cc}
-2&-1\\[0.3cm]
1 &1 \end{array}\right)\\[0.3cm]
\left|\begin{array}{cc}
-2-\lambda&-1\\[0.3cm]
1 &1-\lambda \end{array}\right|=0-\text{characteristic equation}\\[0.3cm]
(-2-\lambda)(1-\lambda)-1\cdot(-1)=\\[0.3cm]
-2+2\lambda-\lambda+\lambda^2+1=0\\[0.3cm]
\lambda^2+\lambda-1=0\\[0.3cm]
D=1^2-4\cdot1\cdot(-1)=1+4=5\to\sqrt{D}=\sqrt{5}\\[0.3cm]
\left[\begin{array}{l}
\lambda_1=\displaystyle\frac{-1+\sqrt{5}}{2}>0,\lambda_1\in\mathbb{R}\\[0.3cm]
\lambda_2=\displaystyle\frac{-1-\sqrt{5}}{2}<0,\lambda_2\in\mathbb{R}\end{array}\right."

Conclusion,


The eigenvalues are real. Eigenvalues opposite sign - An Unstable Saddle Node : rajectories in the general direction of the negative eigenvalue's eigenvector will initially approach the fixed point but will diverge as they approach a region dominated by the positive (unstable) eigenvalue.


ANSWER

(i)



"\boxed{A(0,0)\,\,\,\text{and}\,\,\,B(-1,1)-\text{the real critical points of the system
}}"

(ii)



"\boxed{A(0,0)-\text{An Unstable Spiral}}\\[0.3cm]
\boxed{B(-1,1)-\text{An Unstable Saddle Node}}"